OpenStudy (anonymous):

How do I solve N!/(N-1)!?

zepdrix (zepdrix):

Do you understand what I'm doing in this example here? \[\Large \frac{5!}{4!}\quad=\quad \frac{5\cdot4\cdot3\cdot2\cdot1}{4!}\quad=\quad \frac{5\cdot4!}{4!}\quad=\quad 5\]

OpenStudy (anonymous):

yep

zepdrix (zepdrix):

\[\Large (5-1)!\quad=\quad 4!\] So we could rewrite our example as:\[\Large \frac{5!}{4!}\quad=\quad \frac{5!}{(5-1)!}\]

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

an interesting development...

zepdrix (zepdrix):

lol :)

OpenStudy (anonymous):

so

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

I now the 1x2x3... in both numerators and denominators are canceled out

OpenStudy (anonymous):

I have this

OpenStudy (anonymous):

\[1\times 2 \times 3...(n)/1 \times 2 \times 3...(n-1)\]

zepdrix (zepdrix):

So when n=3,\[\Large \frac{n}{(n-1)!}\quad=\quad \frac{3\cdot2\cdot1}{2\cdot1}\quad=\quad 3\] When n=4,\[\Large \frac{n}{(n-1)!}\quad=\quad \frac{4\cdot3\cdot2\cdot1}{3\cdot2\cdot1}\quad=\quad 4\]What number are we ending up with in each case?

zepdrix (zepdrix):

woops those numerators should be n!, sorry.

OpenStudy (anonymous):

n

zepdrix (zepdrix):

Yesss good good.

OpenStudy (anonymous):

so how would I write an answer out in variables?

OpenStudy (anonymous):

(n)(n-1)/(n-1)?

zepdrix (zepdrix):

Here is another way to think about it. If we take the first number out of that factorial in the numerator (similar to what we did n the example, we're left with this: \[\Large \frac{5\cdot(5-1)!}{(5-1)!}\]

zepdrix (zepdrix):

Ah sorry lemme throw the n back in.

zepdrix (zepdrix):

\[\Large \frac{n!}{(n-1)!}\quad=\quad \frac{n\cdot(n-1)!}{(n-1)!}\]

zepdrix (zepdrix):

Then we have a nice cancellation from there, yes? :o

OpenStudy (anonymous):

but how do we get the (n-1) ontop?

OpenStudy (anonymous):

(n-1) in the numerator that is?

OpenStudy (anonymous):

like, for a problem like (n-1)!/(n-3)!, would the answer be (n-1)(n-3)/(n-3)?

OpenStudy (anonymous):

sorry for all the questions, I hope I am not bugging you

zepdrix (zepdrix):

\[\Large \frac{(n-1)!}{(n-3)!}\]In this example, let's try to play around with the numerator because it starts from a larger number. Look at this example a sec:\[\Large 4!\]If I wanted to pull the 4 out of the factorial, the what does it leave us with in factorial form? The next number below 4, yes?\[\Large 4!\quad=\quad 4\cdot 3!\]Does that setup make sense? :o I just pulled a 4 outside of the factorial thing.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

n=4 for example, would be (4-1)!=3!=3 times 2 times 1=6 in the numberator

OpenStudy (anonymous):

(4-3)!=1!=1 in the denominator?

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

6/1=6

zepdrix (zepdrix):

So what we want to do is, we want to peel numbers off of the larger factorial, until it matches the smaller one.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

so the answer wouldn't be 6n?

zepdrix (zepdrix):

Mmm no not quite :) Stay with me a sec!

zepdrix (zepdrix):

\[\Large \frac{(n-1)!}{(n-3)!}\quad=\quad \frac{(n-1)\cdot(?)!}{(n-3)!}\]So if we pulled the (n-1) out of the top factorial, what number should our factorial start from then? It will be the next number below n-1, which is? :D

OpenStudy (anonymous):

(n-3) :D!!!!!!!!!!!!!!

zepdrix (zepdrix):

No no not quite. that is the number `2 below` n-1. You jumped too far. To get to the number `one below` n-1 we simply subtract 1 ok? So it gives us n-2

zepdrix (zepdrix):

\[\Large \frac{(n-1)!}{(n-3)!}\quad=\quad \frac{(n-1)\cdot(n-2)!}{(n-3)!}\]

OpenStudy (anonymous):

:(, but at least I tried :D

zepdrix (zepdrix):

XD

zepdrix (zepdrix):

So now we want to deal with our new factorial on top, it's still too large. We'll peel another number out of it.\[\Large \frac{(n-1)\cdot(n-2)!}{(n-3)!}\quad=\quad \frac{(n-1)\cdot(n-2)\cdot(?)!}{(n-3)!}\]So we pulled the largest number (n-2) out of the factorial. What will our new factorial start from? It would be the number just below n-2, which is? :O

OpenStudy (anonymous):

(n-3) :DDD!!!!!!

zepdrix (zepdrix):

Good good, that's (n-2)-1, just one number below n-2!

OpenStudy (anonymous):

I think I got the answer

zepdrix (zepdrix):

\[\Large =\quad \frac{(n-1)\cdot(n-2)\cdot\color{#CC0033}{(n-3)!}}{\color{#CC0033}{(n-3)!}}\]

OpenStudy (anonymous):

n squared -3n+2

OpenStudy (anonymous):

:D?

zepdrix (zepdrix):

Yesss good job! :)

OpenStudy (anonymous):

thank you so much

zepdrix (zepdrix):

These problems can be a little tricky. You start with two factorials and you need to try and match them up by shrinking down the larger one.

zepdrix (zepdrix):

\c:/ np